I need the formula to calculate a loan balance as payments are made.
Principal = 150,000
Interest rate = 8.75
Term = 30 years

Answer by rfer(16322)

R=(Pi/12)/(1-(1+i/12)^-n)
P=principal
i=interest rate per pay period per yr.
n=number of payments
R=regular payment
R=(150000*0.00729)/1-(1.00729)^-360
R=1093.5/1-0.0732
R=1093.5/0.9268
R=$1,179.87